OpenStudy (anonymous):

Which of the following shows the square completed in the this equation? x^2 + 3x - 5 = 0 A.(x + 9)2 = 14 B.(x - 3)2 = 8 C.(x + 1.5)2 = 7.25

OpenStudy (anonymous):

is it x^2

OpenStudy (anonymous):

hello???

OpenStudy (anonymous):

Yes i fixed it.

OpenStudy (anonymous):

answer is (x+1.5)^2=7.25

OpenStudy (anonymous):

do u want explanation ??

OpenStudy (anonymous):

(x)^2+2(x)(1.5)+(1.5)^2=5+(1.5)^2